OpenStudy (anonymous):

Factor the trinomial completely x^2-9x+20

OpenStudy (anonymous):

(x-4)(x-5) ? i think

OpenStudy (anonymous):

(x-4)*(x-5) x*x=x^2 -4*x +(-5 *x) = -9x -5* -4= 20